OpenStudy (anonymous):

Pls help!!!!!! how do we find the inverse of below function in terms of f inverse r(x)=1-2f(3-4x)

OpenStudy (anonymous):

we can do this

OpenStudy (anonymous):

start as usual by writing \[y=1-2f(3-4x)\] then switch x and y to get \[x=1-2f(3-4y)\] and solve for \(y\)

OpenStudy (anonymous):

subtract 1 \[x-1=-2f(3-4y)\] divide by -2 \[\frac{1-x}{2}=f(3-4y)\]

OpenStudy (anonymous):

now comes the \(f^{-1}\) part since we want to solve for what is inside the function \[f^{-1}(\frac{1-x}{2})=3-4y\] subtract 3 \[f^{-1}(\frac{1-x}{2})-3=-4y\] divide by -4 \[\frac{3-f^{-1}(\frac{1-x}{2})}{4}=y\] and we are done

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw